determine the horizontal force P required to start moving the 500N block up the inclined surface.assume static friction is 0.3 and angle is 30
1
Expert's answer
2020-09-11T09:15:22-0400


On this picture "\mathbf{mg}" denotes the gravity force, "\mathbf{F_{fric}} = \mu\mathbf{N}" - static frictional force and "\mathbf{N}" - normal force. According to the problem, "mg = 500 N", "\mu = 0.3" and "\theta = 30\degree".

According to the Newton's second law, the projections on the y axis will be:


"N - mg\cos\theta - P\sin\theta = 0\\
N =mg\cos\theta + P\sin\theta"


The projections on the x axis will be:


"F_{fric} + mg\sin\theta - P\cos\theta = 0\\
\mu N + mg\sin\theta - P\cos\theta = 0\\"

Substituting "N" from the first expression, get:


"\mu (mg\cos\theta + P\sin\theta) + mg\sin\theta - P\cos\theta = 0"

Expressing "P", obtain:


"P = mg\dfrac{\mu + \tan\theta}{1 + \mu\tan\theta}\\
\space\\
P = 500\cdot \dfrac{0.3 + \tan30\degree}{1 + 0.3\cdot \tan30\degree} \approx 373.9 N"

Answer. 373.9 N.